OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (1, –7); Slope: A. y-7=-⅔(x+1) B. y-7=⅔(x-1) C. y+7=⅔(x+1) D. y+7=-⅔(x-1)

OpenStudy (mathmale):

the GENERAL point-slope formula is \[y-y _{1}=m(x-x _{1})\]

OpenStudy (mathmale):

You are given the point (1,-7) , but you have not specified what the slope, m, is. Do that now, please.

OpenStudy (anonymous):

oh sorry \[-\frac{ 2 }{ 3 }\]

OpenStudy (texaschic101):

basically, this is just a matter of subbing... y - y1 = m(x - x1) (1,-7)....x1 = 1 and y1 = -7 slope(m) = -2/3 now try subbing into the formula and what do you get ?

OpenStudy (anonymous):

is it d?

OpenStudy (texaschic101):

let me start you off... y - y1....so what is y1...it is -7.....so your problem now reads y - (-7).. = m.......= 2/3.....(x - x1).......(x - 1) so it now reads : y - (-7) = 2/3(x - 1)....but the negatives....so when simplified is : y + 7 = 2/3(x - 1).....it is D

OpenStudy (anonymous):

Oh

OpenStudy (mathmale):

Ashley, it's really important that you show your work if you want meaningful feedback. Please don't respond with "Is it d?" That sounds as tho' y ou're guessing.

OpenStudy (anonymous):

sorry

OpenStudy (mathmale):

texaschick has done a great job of demonstrating what you could do to communicate your work to potential helpers. Thanks, t-c.

OpenStudy (texaschic101):

I was trying not to confuse her too much...I hope she understands the problem..

OpenStudy (mathmale):

And what does Ashley herself think?

OpenStudy (anonymous):

I do understand it now.

OpenStudy (texaschic101):

That's always good to hear....you just have to be sure and pay attention to the signs

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Thank you for helping